need help urgent!
1. A water tank has two valves, A and B. When only valve A is open, the tank is emptied in 6 hours. Furthermore, it takes twelves hours to empty the tank if only valve B is open. How long will it take to empty the tank if both valves are open?
2. An 8-liter solution is 25% alcohol. How much water must be added to reduce the concentration of alcohol in the solution to 20%?

Answer by alicealc(167)

Number 1.
time needed if both valve are opened = T
1/T = 1/A + 1/B
1/T = 1/6 + 1/12 = 2/12 + 1/12 = 3/12 = 1/4
T = 4 hours

Number 2.
Volume of water added = V
25/100 * 8 = 20/100 * (8 + V)
2 = 0.2*(8 + V)
2 = 1.6 + 0.2V
2 - 1.6 = 0.2V
0.4 = 0.2V
V = 0.4/0.2 = 2 liters